OpenStudy (anonymous):

Coach Johnson is 10 years older than Tim. When Tim started football 3 years ago, Coach Johnson was twice his age. How old is Tim now? Numerical Answers Expected!

OpenStudy (anonymous):

John = J T = Tim J - 10 = T T - 3 = J / 2 J = (T - 3) * 2

OpenStudy (anonymous):

x = tim y = coach x+10=y x-3=y/2 2x-6=y using simultenous to both equations y=x+10 y=2x-6 u'll get 16 and 26 :)

OpenStudy (anonymous):

x+10=y is referring to know, which coach is 10 years older 2x-6=y is referring to 3 years ago, when coach is twice age as tim :D

OpenStudy (anonymous):

is referring to now* sorry for typo

OpenStudy (anonymous):

do ask if u have question :)

OpenStudy (anonymous):

No I Under Stand